# Optimal Non-Adaptive Group Testing with One-Sided Error Guarantees

### v1

@@ I-A Problem Setup
Given a set of items $[n]\coloneqq\{1,2,\dots n\}$, there exists some set $S\subseteq[n]$ of defective items with $\lvert S\rvert=k$, where $k=\Theta(n^{\theta})$ for some $\theta\in(0,1)$ and $k$ is known a priori. The set of all possible values of $S$ is denoted by ${\mathcal{S}}$, and $S$ is assumed to be uniformly distributed over all ($n\choose k$ elements in ${\mathcal{S}}$, which is known as the combinatorial prior. We also consider in certain parts of our proofs the i.i.d prior (e.g., as an intermediate step), where each item is defective independently with some probability $q\in(0,1)$.
@@
For the SUBSET and SUPERSET problems, the error probabilities are given as follows with parameters $\eta^{-},\eta^{+}>0$:
@@ I-B Prior Work
In the sparse setting (namely, $k=\Theta(n^{\theta})$ with $\theta\in(0,1)$), there exists a ubiquitous lower bound of $\big{(}T\geq k\log_{2}\frac{n}{k}\big{)}(1-o(1))$ on the number of tests required to obtain a vanishing probability of error, regardless of adaptivity or non-adaptivity. This bound is known as the counting bound [13]. The intuition behind this bound stems from the fact that $\log_{2}{n\choose k}=\big{(}k\log_{2}\frac{n}{k}\big{)}\big{(}1+o(1)\big{)}$ bits are required to describe the defective set, and each test can provide at most 1 bit of information since the test outcomes are binary. This intuition can easily be made precise using standard information theoretic tools such as Fano’s inequality [14, Theorem 1], which implies we cannot obtain an arbitrarily small probability of error with fewer tests. This result has also been strengthened to obtain a strong converse bound stating that $\mathbb{P}(\hat{S}\neq S)\to 1$ as $n\to\infty$ when $T\leq\left(k\log_{2}\frac{n}{k}\right)(1-\varepsilon)$ for arbitrarily small $\varepsilon>0$ [15, Theorem 1].
The counting bound can be met when adaptive testing is used for all $\theta\in(0,1)$ using a generalized binary splitting algorithm due to Hwang [16]. The counting bound can also be achieved for $\theta\in\big{(}0,\frac{\log 2}{1+\log 2}\big{)}\approx(0,0.409)$ non-adaptively, while a converse bound specific to the non-adaptive setting shows that more tests are required for $\theta\geq\frac{\log 2}{1+\log 2}$ [17]. Furthermore, [17, Theorem 2] shows that there exists a polynomial-time algorithm matching this bound in the non-adaptive setting. (Prior to their work, the bound was showed to be information-theoretically achievable in [18].)
Approximate recovery. There is also a converse bound specific to the approximate recovery setting [15, Theorem 3], which states that the high probability recovery of a set $\hat{S}$ such that $\max\big{\{}\lvert\hat{S}\setminus S\rvert,\lvert S\setminus\hat{S}\rvert\big{\}}\leq\beta k$ for fixed $\beta\in(0,1)$ is not possible when the number of tests satisfies $T\leq(1-\beta)\left(k\log_{2}\frac{n}{k}\right)(1-\varepsilon)$ for arbitrarily small $\varepsilon>0$. When an equal number of false positive and false negative items are allowed (i.e., two-sided approximate recovery), a matching upper bound can be achieved [12]. The polynomial-time algorithm from [17] can also be used to achieve the same bound.
@@
There are also some related works tackling similar problems to SUBSET and SUPERSET. Aldridge [20] studied a related problem of isolating defectives, and translated to our setup, one of their results solves an instance of the SUBSET problem using a combination of the exact recovery algorithm in [17] and a proposed algorithm using ideas from SAFFRON [21]. However, in the regimes where the latter algorithm attains the better rate, it only recovers roughly $58\%$ of the defective set and attains a rate below $0.19$, which turns out to be highly suboptimal. Sharma et al. [22] also considered the problem of finding a subset of healthy individuals in the population, but taking the complement of that set would give a “very enlarged superset” rather than “slightly enlarged” as we seek. The SUPERSET problem is also related to list decoding, which has been considered in works such as [23, 24] but with very different goals to ours (e.g., as a stepping stone to exact recovery). Finally, Mézard et al. [25] established a converse bound stating that a rate above $\log 2$ is not achievable for a problem related to SUPERSET. However, their result concerns two-stage adaptive procedures rather than the non-adaptive designs we consider, and they are interested in the zero-error recovery problem rather than the small-error criterion that we consider.
@@ I-C Achievable Rates Adapted from Prior Works
Then if $\eta^{-}=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ for any $\lambda\in[0,1)$, and $\eta^{+}=k^{-o(1)}$, the rate $R^{*}$ is achievable in both the SUBSET and SUPERSET problems.
The condition $\eta^{+}=k^{-o(1)}$ covers a range of possible values, including not only $\eta^{+}=\Theta(1)$ but also the case of decaying to zero sufficiently slowly (e.g., as $\frac{1}{(\log k)^{c}}$ for some $c>0$). Moreover, the condition on $\eta^{-}$ allows for even faster (polynomial) decay to zero; by letting $\lambda$ be close to $1$ we can allow $\eta^{-}=\big{(}\frac{k}{n}\big{)}^{1-\varepsilon}$ for arbitrarily small $\varepsilon>0$, amount to at most $k\big{(}\frac{k}{n}\big{)}^{1-\varepsilon}$ false negatives in the SUBSET problem. Note also that the term $\zeta(\theta)$ is the optimal exact recovery rate (among all designs) from [17], and is present because achieving exact recovery implies solving both problems even with $\eta^{+},\eta^{-}=0$. This term is dominant for $\theta<\frac{1}{2}$, and for any such $\theta$ the quantity $\eta^{-}k=\Theta(k(\frac{k}{n})^{\lambda})$ is $o(1)$ for $\lambda$ sufficiently close to one, which means that the “approximate recovery” guarantee of SUBSET is in fact equivalent to exact recovery.
@@ I-D Contributions
  • We provide a new algorithm for the SUBSET problem that achieves a rate of 1 for all $\theta\in(0,1)$ when $\eta^{-}=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ for any $\lambda\in(0,\frac{1}{2})$, thus matching the counting bound and beating the best existing achievability result.
  • We provide a strong converse bound for the SUPERSET problem that shows the rate in (5) is optimal even for $\eta^{+}=k^{o(1)}$ among arbitrary non-adaptive designs, with this converse holding for even larger $\eta^{+}$ when $\theta\in(0.409,\frac{1}{2})$, namely, $\eta^{+}=O(k^{\lambda})$ with $\lambda\in(0,\frac{1}{\theta}-2]$.
@@ I-E Notation
We generally use bold face letters to denote vectors, lowercase to denote scalars and capital letters to denote random variables (though the number of tests $T$ is an exception). We use standard Bachmann-Landau notation $o(\cdot),\ \omega(\cdot),\ O(\cdot),\ \Omega(\cdot),\ \Theta(\cdot)$ to denote asymptotic relations. For a matrix ${\mathsf{M}}$ (resp. vector $\mathbf{v}$) and set ${\mathcal{J}}\subseteq[n]$, ${\mathsf{M}}_{{\mathcal{J}}}$ denotes the sub-matrix of ${\mathsf{M}}$ (resp. sub-vector of $\mathbf{v}$) corresponding to the columns (resp. entries) indexed by ${\mathcal{J}}$. If ${\mathcal{J}}=\emptyset$ we use the convention that ${\mathsf{M}}_{\emptyset}$ (resp. $\mathbf{v}_{\emptyset}$) =absent=$\emptyset$= ∅. For a set ${\mathcal{N}}$ and some ${\mathcal{U}}\subseteq{\mathcal{N}}$, we denote ${\mathcal{U}}^{\mathsf{c}}\coloneqq{\mathcal{N}}\setminus{\mathcal{U}}$ as the complement of ${\mathcal{U}}$. All logarithms have base $e$ unless specified otherwise (e.g., via $\log_{2}$).
The Hamming distance between two vectors $\mathbf{u},\mathbf{v}\in\{0,1\}^{n}$ is given by $d_{\mathrm{H}}(\mathbf{u},\mathbf{v})$. We also slightly abuse notation throughout and use $d_{\mathrm{H}}(S,S^{\prime})$ for $S,S^{\prime}\subseteq[n]$, which is to be understood as the Hamming distance between the binary vectors having $1$’s in the locations indexed by $S,S^{\prime}$ and 00’s elsewhere.
@@ II Main Results
In the SUBSET problem, for any $\theta\in(0,1)$, there exists an algorithm that achieves a rate of 1 for recovering an estimate $\hat{S}$ of size $(1-\eta^{-})k$, when $\eta^{-}=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ for any $\lambda\in(0,\frac{1}{2})$.
@@
cannot be achieved regardless of test design. In particular, for rates above $R^{*}$, it holds that $P_{\mathrm{e}}^{+}\to 1$. Moreover, when $\theta\in\big{(}\frac{\log 2}{1+\log 2},\frac{1}{2}\big{)}$ the same result also holds for $\eta^{+}=O(k^{\lambda})$, for any $\lambda\in(0,\frac{1}{\theta}-2]$.
@@
For $\varepsilon>0$ and any fixed $\alpha\in(0,1)$, there exists an algorithm that outputs an estimate $\hat{S}$ of $S$ of size $(1-\alpha)k$ such that $\hat{S}\subseteq S$ with probability $1-o(1)$ whenever $T\geq(1-\alpha)(k\log_{2}\frac{n}{k})(1+\varepsilon)$. In other words, a rate of $\frac{1}{1-\alpha}$ is achievable.
@@
While we attempted to find the best possible range of $\lambda$ in Theorems 2 and 3 based on what our own analysis can give, we do not claim these to be the best possible among all designs. In particular, we expect that adopting a near constant column weight design for SUBSET would lead to higher range of $\lambda$ being allowed. [^1] One way of seeing this is to observe that setting $\theta<\frac{1}{3}$ and $\lambda\approx\frac{1}{2}$ in Theorem 2 permits $\eta^{-}<\frac{1}{k}$, which is equivalent to exact recovery. This is consistent with exact recovery being possible for all $\theta\leq\frac{1}{3}$ under the Bernoulli design. For the near-constant column weight design, this range increases to $\theta\leq\frac{\log 2}{1+\log 2}\approx 0.409$, which suggests that the range of allowed $\lambda$ should be higher, namely $\lambda\in(0,\log 2)$. We stick to the Bernoulli design since it simplifies the analysis, and because making $\lambda$ as high as possible is not our main focus.
@@ III-A Proof of Theorem 2 (Achievability for SUBSET)
where $\gamma_{\lfloor\beta k+1\rfloor},\gamma_{\lfloor\beta k+2\rfloor},\dots,\gamma_{k}$ are positive constants and $\imath^{T}\left({\mathsf{X}}_{s_{\mathrm{dif}}};\mathbf{y}\mid{\mathsf{X}}_{s_{\mathrm{eq}}}\right)$ is a conditional information density (the precise definition is omitted as it is not important for our purposes). The proof of [15, Theorem 3] reveals that for appropriate choices of $\gamma_{\lfloor\beta k+1\rfloor},\gamma_{\lfloor\beta k+2\rfloor},\dots,\gamma_{k}$ the algorithm succeeds in finding such a set if $\beta\in(0,1)$ is fixed and $T\geq\left(k\log_{2}\frac{n}{k}\right)(1+\varepsilon)$ for arbitrarily small $\varepsilon>0$. It turns out that their analysis can also be extended from constant $\beta$ to any $\beta=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ with $\lambda\in(0,\frac{1}{2})$, with only minor changes. The proof of this slightly strengthened statement is outlined in Appendix -B.
@@
Input: Estimate $\hat{S}^{\prime}$ from the approximate recovery algorithm [15] with $\beta=\frac{\eta^{-}}{2}$, test matrix ${\mathsf{X}}$, test outcome vector $\mathbf{y}$
@@
The SUBSET algorithm first takes the estimate returned from the approximate recovery algorithm (which has size $k$), and then searches through all subsets of size $(1-\eta^{-})k$ within Hamming distance at most $2\eta^{-}k$ from the approximate recovery estimate, returning the set that explains the most tests. Note that this algorithm does not conduct any additional tests outside of the ones conducted to obtain $\hat{S}^{\prime}$. As such, the testing matrix ${\mathsf{X}}$ follows a Bernoulli test design with the following inclusion probability:
$p=\frac{\log 2}{k}\big{(}1+o(1)\big{)}\leavevmode\nobreak\ \text{ satisfying }\leavevmode\nobreak\ \mathbb{P}(Y_{t}=1)=\frac{1}{2}.$ (11)
The algorithm is successful if the returned set $\hat{S}$ is a subset of the defective set $S$. We will prove that this is indeed the case with high probability. To show this, it suffices to show that a subset of $S$ explains the most tests with high probability among all sets of size $(1-\eta^{-})k$ within distance $2\eta^{-}k$ of the approximate recovery estimate $\hat{S}^{\prime}$. To this end, a first thought would be to condition on the approximate recovery estimate $\hat{S}^{\prime}$ and analyze the sets around this estimate. However, this approach appears to be difficult, since conditioning on this event means that the entries of ${\mathsf{X}}$ are no longer independent, which complicates the analysis.
@@
We know that $d_{\mathrm{H}}(\hat{S}^{\prime},S)\leq\eta^{-}k$ (with high probability) due to the choice $\beta=\frac{\eta^{-}}{2}$ in Algorithm 1, and when this holds, the triangle inequality gives ${\mathcal{S}}_{\mathrm{alg}}\subseteq{\mathcal{S}}_{\mathrm{an}}$. In other words, if we prove a certain property for ${\mathcal{S}}_{\mathrm{an}}$, we can transfer that property to ${\mathcal{S}}_{\mathrm{alg}}$. When studying ${\mathcal{S}}_{\mathrm{an}}$, there is no need for conditioning of the kind discussed above.
@@
If $T=\big{(}k\log_{2}\frac{n}{k}\big{)}(1+\varepsilon)$ for some $\varepsilon>0$, and $\eta^{-}=\Omega((\frac{k}{n})^{\lambda})$ for some $\lambda\in\big{(}0,\frac{1}{2}\big{)}$, [^2] We note that $k^{-\lambda(\frac{1}{\theta}-1)}=\Theta((\frac{k}{n})^{\lambda})$; the former parametrization is mainly used for the theorem statement while the latter is more convenient for this and subsequent proofs. then for any $\delta_{1}=1-o(1)$, it holds with probability $1-o(1)$ that there are at least $k-\frac{\eta^{-}}{2}k$ defective items $i$ such that $G_{i}\geq\frac{1-\delta_{1}}{2}\log\frac{n}{k}$.
@@
where (18) follows since $(1-\delta_{1})\log(1-\delta_{1})+\delta_{1}=1-o(1)$ when $\delta_{1}=1-o(1)$. Let $Z=\sum_{i=1}^{k}\mathds{1}\big{\{}G_{i}\leq\frac{1-\delta_{1}}{2}\log\frac{n}{k}\big{\}}$; by the linearity of expectation, $\mathbb{E}Z\leq k\left(\frac{k}{n}\right)^{\frac{1}{2}-o(1)}$, and by Markov’s inequality, it follows that
$\mathbb{P}\left(Z\geq\frac{\eta^{-}}{2}k\right)$ $\leq\frac{\mathbb{E}Z}{\frac{\eta^{-}}{2}k}$
$\leq\frac{2\left(\frac{k}{n}\right)^{\frac{1}{2}-o(1)}}{\eta^{-}}$
$\leq 2c\left(\frac{k}{n}\right)^{\frac{1}{2}-\lambda-o(1)}$ (19)
@@
where $c$ is the value such that $\frac{1}{\eta^{-}}\leq c(\frac{n}{k})^{\lambda}$ for sufficiently large $n$ (since $\eta^{-}=\Omega((\frac{k}{n})^{\lambda})$), and (20) follows since $\lambda\in(0,\frac{1}{2})$ and $k=o(n)$. Hence, taking the contrapositive, we have with probability $1-o(1)$ that at least $k-\frac{\eta^{-}}{2}k$ defective items $i\in S$ have $G_{i}\geq\frac{1-\delta_{1}}{2}\log\frac{n}{k}$ as claimed. ∎
Lemma 4 implies that the following holds with probability $1-o(1)$:
$\exists\tilde{S}\,:\,\big{(}\lvert\tilde{S}\rvert=(1-\eta^{-})k\big{)}\wedge\big{(}\tilde{S}\subseteq S\big{)}\wedge\bigg{(}\text{each $i\in\tilde{S}$ is in at least $\frac{1-\delta_{1}}{2}\log\frac{n}{k}$ good tests}\bigg{)},$ (21)
@@
which follows from the definition of ${\mathcal{S}}_{\mathrm{alg}}$ in (13) along with $d_{\mathrm{H}}(\tilde{S},S)=|S\setminus\tilde{S}|=\eta^{-}k$, $d_{\mathrm{H}}(\hat{S}^{\prime},S)\leq\eta^{-}k$, and the triangle inequality. The importance of $\tilde{S}$ is highlighted in the following lemma.

### v3

@@ I-A Problem Setup
Given a set of items $n]\coloneqq\{1,2,\dots n\}$, there exists some set $S\subseteq[n]$ of defective items with $\lvert S\rvert=k$, where $k=\Theta(n^{\theta})$ for some $\theta\in(0,1)$ and $k$ is known a priori. The set of all possible values of $S$ is denoted by $\mathcal{S}$}, and $S$ is assumed to be uniformly distributed over all $n\choose k$ elements in $\mathcal{S}$}, which is known as the combinatorial prior. We also consider in certain parts of our proofs the i.i.d prior (e.g., as an intermediate step), where each item is defective independently with some probability $q\in(0,1)$.
@@
For the SUBSET and SUPERSET problems, the error probabilities are given as follows with parameters $\eta^{-},\eta^{+}>0$.
@@
Note that the probabilities in these definitions are over the random defective set $S$ and any potential randomness in the test design and decoder. Moreover, we assume for notational convenience that $(1-\eta^{-})k$ and $(1+\eta^{+})k$ are integers, but if not, all of our results will hold regardless of whether these values are rounded up or rounded down.
@@ I-B Prior Work
In the sparse setting (namely, $k=\Theta(n^{\theta})$ with $\theta\in(0,1)$), there exists a ubiquitous lower bound of $T\geq\big(k\log_{2}\frac{n}{k}\big)(1-o(1))$ on the number of tests required to obtain a vanishing probability of error, regardless of adaptivity or non-adaptivity. This bound is known as the counting bound [13]. The intuition behind this bound stems from the fact that $\log_{2}{n\choose k}=\big(k\log_{2}\frac{n}{k}\big)\big(1+o(1)\big)$ bits are required to describe the defective set, and each test can provide at most 1 bit of information since the test outcomes are binary. This intuition can easily be made precise using standard information theoretic tools such as Fano’s inequality [14, Theorem 1], which implies we cannot obtain an arbitrarily small probability of error with fewer tests. This result has also been strengthened to obtain a strong converse bound stating that $\mathbb{P}(\hat{S}\neq S)\to 1$ as $n\to\infty$ when $T\leq\left(k\log_{2}\frac{n}{k}\right)(1-\varepsilon)$ for arbitrarily small $\varepsilon>0$ [13, Thm. 3.1].
The counting bound can be met when adaptive testing is used for all $\theta\in(0,1)$ using a generalized binary splitting algorithm due to Hwang [15]. The counting bound can also be achieved for $\theta\in\big(0,\frac{\log 2}{1+\log 2}\big)\approx(0,0.409)$ non-adaptively, while a converse bound specific to the non-adaptive setting shows that more tests are required for $\theta\geq\frac{\log 2}{1+\log 2}$ [16]. Furthermore, [16, Theorem 2] shows that there exists a polynomial-time algorithm matching this bound in the non-adaptive setting. (Prior to their work, the bound was showed to be information-theoretically achievable in [17].)
Approximate recovery. There is also a converse bound specific to the approximate recovery setting [18, Theorem 3], which states that the high probability recovery of a set $\hat{S}$ such that $\max\big\{\lvert\hat{S}\setminus S\rvert,\lvert S\setminus\hat{S}\rvert\big\}\leq\beta k$ for fixed $\beta\in(0,1)$ is not possible when the number of tests satisfies $T\leq(1-\beta)\left(k\log_{2}\frac{n}{k}\right)(1-\varepsilon)$ for arbitrarily small $\varepsilon>0$. When an equal number of false positive and false negative items are allowed (i.e., two-sided approximate recovery), a matching upper bound can be achieved [12]. The polynomial-time algorithm from [16] can also be used to achieve the same bound.
@@
There are also some related works tackling similar problems to SUBSET and SUPERSET. Aldridge [20] studied a related problem of isolating defectives, and translated to our setup, one of their results solves an instance of the SUBSET problem using a combination of the exact recovery algorithm in [16] and a proposed algorithm using ideas from SAFFRON [21]. However, in the regimes where the latter algorithm attains the better rate, it only recovers roughly $58\%$ of the defective set and attains a rate below $0.19$, which turns out to be highly suboptimal. Sharma et al. [22] also considered the problem of finding a subset of healthy individuals in the population, but taking the complement of that set would give a “very enlarged superset” rather than “slightly enlarged” as we seek. The SUPERSET problem is also related to list decoding, which has been considered in works such as [23, 24, 10, 25] but with very different goals to ours (e.g., as a stepping stone to exact recovery in the zero-error setting). Mézard et al. [26] established a converse bound stating that a rate above $\log 2$ is not achievable for a problem related to SUPERSET. However, their result concerns two-stage adaptive procedures rather than the non-adaptive designs we consider, and they are interested in the zero-error recovery problem rather than the small-error criterion that we consider.
@@ I-C Achievable Rates Adapted from Prior Works
Then if $\eta^{-}=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ for any $\lambda\in[0,1)$, and $\eta^{+}=k^{o(1)}$, the rate $R^{*}$ is achievable in both the SUBSET and SUPERSET problems.
The condition $\eta^{+}=k^{o(1)}$ covers a range of possible values, including not only $\eta^{+}=\Theta(1)$ but also the case of growing sufficiently slowly (e.g., as $(\log k)^{c}$ for some $c>0$). Moreover, the condition on $\eta^{-}$ allows not only $\eta^{-}=\Theta(1)$ and “slow” decay to zero at rate $k^{-o(1)}$, but also faster (polynomial) decay; by letting $\lambda$ be close to $1$ we can allow $\eta^{-}=\big(\frac{k}{n}\big)^{1-\varepsilon}$ for arbitrarily small $\varepsilon>0$, amounting to at most $k\big(\frac{k}{n}\big)^{1-\varepsilon}$ false negatives in the SUBSET problem. Note also that the term $\zeta(\theta)$ is the optimal exact recovery rate (among all designs) from [16], and is present because achieving exact recovery implies solving both problems even with $\eta^{+},\eta^{-}=0$. This term is dominant for $\theta<\frac{1}{2}$, and for any such $\theta$ the quantity $\eta^{-}k=\Theta(k(\frac{k}{n})^{\lambda})$ is $o(1)$ for $\lambda$ sufficiently close to one, which means that the “approximate recovery” guarantee of SUBSET is in fact equivalent to exact recovery.
@@ I-D Contributions
  • We provide a new algorithm for the SUBSET problem that achieves a rate of 1 for all $\theta\in(0,1)$ when $\eta^{-}=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ for any $\lambda\in[0,\frac{1}{2})$, thus matching the counting bound and beating the best existing achievability result.
  • We provide a strong converse bound for the SUPERSET problem that shows the rate in (5) is optimal for any $\eta^{+}=k^{o(1)}$ among arbitrary non-adaptive designs, with this converse holding for even larger $\eta^{+}$ when $\theta\in(0.409,\frac{1}{2})$, namely, $\eta^{+}=O(k^{\lambda})$ with $\lambda\in[0,\frac{1}{\theta}-2]$.
@@ I-E Notation
We generally use bold face letters to denote vectors, lowercase to denote scalars and capital letters to denote random variables (though the number of tests $T$ is an exception). We use standard Bachmann-Landau notation $o(\cdot),\ \omega(\cdot),\ O(\cdot),\ \Omega(\cdot),\ \Theta(\cdot)$ to denote asymptotic relations. For a matrix $\mathsf{M}$} (resp. vector $\mathbf{v}$) and set ${\mathcal{J}}\subseteq[n]$, ${\mathsf{M}}_{{\mathcal{J}}}$ denotes the sub-matrix of $\mathsf{M}$} (resp. sub-vector of $\mathbf{v}$) corresponding to the columns (resp. entries) indexed by $\mathcal{J}$}. If $\mathcal{J}$$\emptyset$ we use the convention that $\mathsf{M}$$_{\emptyset}$ (resp. $\mathbf{v}_{\emptyset}$) $\emptyset$. For a set $\mathcal{N}$} and some ${\mathcal{U}}\subseteq{\mathcal{N}}$, we denote ${\mathcal{U}}^{\mathsf{c}}\coloneqq{\mathcal{N}}\setminus{\mathcal{U}}$ as the complement of $\mathcal{U}$}. All logarithms have base $e$ unless specified otherwise (e.g., via $\log_{2}$).
The Hamming distance between two vectors $\mathbf{u},\mathbf{v}\in\{0,1\}^{n}$ is given by $d_{\mathrm{H}}(\mathbf{u},\mathbf{v})$. We also slightly abuse notation throughout and use $d_{\mathrm{H}}(S,S^{\prime})$ for $S,S^{\prime}\subseteq[n]$, which is to be understood as the Hamming distance between the binary vectors having $1$’s in the locations indexed by $S,S^{\prime}$ and 0’s elsewhere.
@@ II Main Results
In the SUBSET problem, for any $\theta\in(0,1)$, there exists an algorithm that achieves a rate of 1 for recovering an estimate $\hat{S}$ of size $(1-\eta^{-})k$, when $\eta^{-}=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ for any $\lambda\in[0,\frac{1}{2})$.
@@
cannot be achieved regardless of test design. In particular, for rates above $R^{*}$, it holds that $P_{\mathrm{e}}^{+}\to 1$. Moreover, when $\theta\in\big(\frac{\log 2}{1+\log 2},\frac{1}{2}\big)$ the same result also holds for $\eta^{+}=O(k^{\lambda})$, for any $\lambda\in[0,\frac{1}{\theta}-2]$.
@@
For $\varepsilon>0$ and any fixed $\alpha\in(0,1)$, there exists an algorithm that outputs an estimate $\hat{S}$ of $S$ of size $(1-\alpha)k$ such that $\hat{S}\subseteq S$ with probability 1o(1)1-o(1) whenever $T\geq(1-\alpha)(k\log_{2}\frac{n}{k})(1+\varepsilon)$. In other words, a rate of $\frac{1}{1-\alpha}$ is achievable.
@@
While we attempted to find the best possible range of $\lambda$ in Theorems 2 and 3 based on what our own analysis can give, we do not claim these to be the best possible among all designs. In particular, we expect that adopting a near constant column weight design for SUBSET would lead to higher range of $\lambda$ being allowed. [^1] One way of seeing this is to observe that setting $\theta<\frac{1}{3}$ and $\lambda\approx\frac{1}{2}$ in Theorem 2 permits $\eta^{-}<\frac{1}{k}$, which is equivalent to exact recovery. This is consistent with exact recovery being possible for all $\theta\leq\frac{1}{3}$ under the Bernoulli design. For the near-constant column weight design, this range increases to $\theta\leq\frac{\log 2}{1+\log 2}\approx 0.409$, which suggests that the range of allowed $\lambda$ should be higher, namely $\lambda\in[0,\log 2)$. We stick to the Bernoulli design since it simplifies the analysis, and because making $\lambda$ as high as possible is not our main focus.
@@ III-A Proof of Theorem 2 (Achievability for SUBSET)
where $\gamma_{\lfloor\beta k+1\rfloor},\gamma_{\lfloor\beta k+2\rfloor},\dots,\gamma_{k}$ are positive constants and $\imath^{T}\left({\mathsf{X}}_{s_{\mathrm{dif}}};\mathbf{y}\mid{\mathsf{X}}_{s_{\mathrm{eq}}}\right)$ is a conditional information density (the precise definition is omitted as it is not important for our purposes). The proof of [18, Theorem 3] reveals that for appropriate choices of $\gamma_{\lfloor\beta k+1\rfloor},\gamma_{\lfloor\beta k+2\rfloor},\dots,\gamma_{k}$ the algorithm succeeds in finding such a set if $\beta\in(0,1)$ is fixed and $T\geq\left(k\log_{2}\frac{n}{k}\right)(1+\varepsilon)$ for arbitrarily small $\varepsilon>0$. It turns out that their analysis can also be extended from constant $\beta$ to any $\beta=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ with $\lambda\in[0,\frac{1}{2})$, with only minor changes. The proof of this slightly strengthened statement is outlined in Appendix -B.
@@
Input: Estimate $\hat{S}^{\prime}$ from the approximate recovery algorithm [18] with $\beta=\eta^{-}$, test matrix $\mathsf{X}$}, test outcome vector $\mathbf{y}$
@@
The SUBSET algorithm first takes the estimate returned from the approximate recovery algorithm (which has size $k$), and then searches through all subsets of size $(1-\eta^{-})k$ within Hamming distance at most $3\eta^{-}k$ from the approximate recovery estimate, returning the set that explains the most tests. Note that this algorithm does not conduct any additional tests outside of the ones conducted to obtain $\hat{S}^{\prime}$. As such, the testing matrix $\mathsf{X}$} follows a Bernoulli test design with the following inclusion probability:
$p=\frac{\log 2}{k}\big(1+o(1)\big)\penalty 10000\ \text{ satisfying }\penalty 10000\ \mathbb{P}(Y_{t}=1)=\frac{1}{2}.$ (11)
The algorithm is successful if the returned set $\hat{S}$ is a subset of the defective set $S$. We will prove that this is indeed the case with high probability. To show this, it suffices to show that a subset of $S$ explains the most tests with high probability among all sets of size $(1-\eta^{-})k$ within distance $3\eta^{-}k$ of the approximate recovery estimate $\hat{S}^{\prime}$. To this end, a first thought would be to condition on the approximate recovery estimate $\hat{S}^{\prime}$ and analyze the sets around this estimate. However, this approach appears to be difficult, since conditioning on this event means that the entries of $\mathsf{X}$} are no longer independent, which complicates the analysis.
@@
We know that $d_{\mathrm{H}}(\hat{S}^{\prime},S)\leq 2\eta^{-}k$ (with high probability) due to the choice $\beta=\eta^{-}$ in Algorithm 1, and when this holds, the triangle inequality gives ${\mathcal{S}}_{\mathrm{alg}}\subseteq{\mathcal{S}}_{\mathrm{an}}$. In other words, if we prove a certain property for ${\mathcal{S}}_{\mathrm{an}}$, we can transfer that property to ${\mathcal{S}}_{\mathrm{alg}}$. When studying ${\mathcal{S}}_{\mathrm{an}}$, there is no need for conditioning of the kind discussed above.
@@
If $T=\big(k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$ for some $\varepsilon>0$, and $\eta^{-}=\Omega((\frac{k}{n})^{\lambda})$ for some $\lambda\in\big[0,\frac{1}{2}\big)$, [^2] We note that $k^{-\lambda(\frac{1}{\theta}-1)}=\Theta((\frac{k}{n})^{\lambda})$; the former parametrization is mainly used for the theorem statement while the latter is more convenient for this and subsequent proofs. then for any $\delta_{1}=1-o(1)$, it holds with probability 1o(1)1-o(1) that there are at least $k(1-\eta^{-})$ defective items $i$ such that $G_{i}\geq\frac{1-\delta_{1}}{2}\log\frac{n}{k}$.
@@
where (18) follows since $(1-\delta_{1})\log(1-\delta_{1})+\delta_{1}=1-o(1)$ when $\delta_{1}=1-o(1)$. Let $Z=\sum_{i\in S}\mathds{1}\big\{G_{i}\leq\frac{1-\delta_{1}}{2}\log\frac{n}{k}\big\}$; by the linearity of expectation, $\mathbb{E}Z\leq k\left(\frac{k}{n}\right)^{\frac{1}{2}-o(1)}$, and by Markov’s inequality, it follows that
$\mathbb{P}\left(Z\geq\eta^{-}k\right)$ $\leq\frac{\mathbb{E}Z}{\eta^{-}k}$
$\leq\frac{\left(\frac{k}{n}\right)^{\frac{1}{2}-o(1)}}{\eta^{-}}$
$\leq c\left(\frac{k}{n}\right)^{\frac{1}{2}-\lambda-o(1)}$ (19)
@@
where $c$ is the value such that $\frac{1}{\eta^{-}}\leq c(\frac{n}{k})^{\lambda}$ for sufficiently large $n$ (since $\eta^{-}=\Omega((\frac{k}{n})^{\lambda})$), and (20) follows since $\lambda\in[0,\frac{1}{2})$ and $k=o(n)$. Hence, taking the contrapositive, we have with probability 1o(1)1-o(1) that at least $k-\eta^{-}k$ defective items $i\in S$ have $G_{i}\geq\frac{1-\delta_{1}}{2}\log\frac{n}{k}$ as claimed. ∎
Lemma 4 implies that the following holds with probability 1o(1)1-o(1):
[…]
@@
which follows from the definition of ${\mathcal{S}}_{\mathrm{alg}}$ in (13) along with $d_{\mathrm{H}}(\tilde{S},S)=|S\setminus\tilde{S}|=\eta^{-}k$, $d_{\mathrm{H}}(\hat{S}^{\prime},S)\leq 2\eta^{-}k$, and the triangle inequality. The importance of $\tilde{S}$ is highlighted in the following lemma.